### v1

@@ Abstract
Scientists across disciplines often use data from the internet to conduct research, generating valuable insights about human behavior. However, as generative AI relying on massive text corpora becomes increasingly valuable, platforms have greatly restricted access to data through official channels. As a result, researchers will likely engage in more web scraping to collect data, introducing new challenges and concerns for researchers. This paper proposes a comprehensive framework for web scraping in social science research for U.S.-based researchers, examining the legal, ethical, institutional, and scientific factors that researchers should consider when scraping the web. We present an overview of the current regulatory environment impacting when and how researchers can access, collect, store, and share data via scraping. We then provide researchers with recommendations to conduct scraping in a scientifically legitimate and ethical manner. We aim to equip researchers with the relevant information to mitigate risks and maximize the impact of their research amidst this evolving data access landscape.
@@ 3.1.1 Contractual restrictions
Researchers should, however, exercise caution before concluding that they are not bound by browsewrap agreements that restrict scraping. For starters, in the Berman v Freedom Financial Network, LLC decision, the court noted that if “the website operator can show that a consumer has actual knowledge of the agreement” the terms prohibiting scraping may be enforceable (Berman v. Freedom Financial Network, LLC, 2022, p. 856). What’s more, cases where courts have set aside anti-scraping provisions have typically involved fact patterns where the defendant scraped only publicly-available information or information that the relevant user was already permitted to access through ordinary use of the service. Courts may be less willing to permit scraping that involves content protected by technical safeguards and/or accessible only to certain classes of users. [^5] In this context, researchers’ deploying other users to assist in scraping – such as through the installation of browser extensions that can scrape within their user accounts – could lead to breach of contract claims, at least with regard to those users who have accepted the platform’s user agreement.
@@ 3.4.1 Poor sampling frames
The key takeaway from these sampling frame issues is that the overall representativeness of research based on scraped data is extremely difficult to confirm. For example Tromble et al. (2017) and González-Bailón et al. (2014) find significant variation in Twitter datasets depending on the sampling frame and API endpoints used. With scraped data, however, researchers face a higher level of uncertainty surrounding the sampling frame and data returned, as there is typically no documentation regarding what researchers can expect to reliability collect, as is the case with APIs. Given this limitation, researchers should carefully consider any claims they make with regard to the broad applicability of their findings.
@@ 4 Recommendations
Different data access methods necessitate different legal, scientific, ethical, and institutional considerations. Importantly, researchers should note that the same data collected in different ways may invoke different implications. In the following section, we make recommendations for a set of best practices that U.S.-based researchers can follow (and adapt to their contexts). In addition to immediate recommendations for how to conduct research using scraped data right now, we also provide recommendations for researchers to improve how we take on data scraping projects in the long term.
@@ 4.3 Institutional
As we highlight in Section 3.3, the institutionalization of the ethical, legal, and technical considerations aims to limit risks for researchers, in part by leveraging the expertise that exists across institutional contexts. A key challenge for researchers is determining who to engage, when, and towards what ends. The Institutional Review Board will ultimately be responsible for ensuring that a project conforms with internal policies and legal standards for research. Put simply, researchers should err on the side of engaging their IRB early and often. An IRB should be able to provide guidance on how it understands the risks associated with scraping, as well as connect researchers to other resources from across the university that can help mitigate those risks. Those resources include IT services, libraries, and general counsel. For researchers engaging in scraping for the first time—or for the first time at a particular institution—colleagues can also act as a resource for understanding how scraping is viewed within that particular environment and how to best navigate the specific actors who are involved in institutionalizing the considerations discussed above.
@@ 4.4 Scientific
First, researchers should make as much effort as possible for the project to develop a sampling strategy—as opposed to simply using convenience as their guide. One laudable example is the way in which The Markup developed a sample for their scraping project, Citizen Browser. Unlike many other studies which relied on snowball or other convenience samples for data collection, The Markup started their study by recruiting a panel of Facebook users based on a survey vendor’s national sample. Only then did they install scraping tools connected to those users accounts, which allowed them to appropriately re-weight their data to allow for population inference (The Citizen Browser Project—Auditing the Algorithms of Disinformation – The Markup, 2020). Because of the rigorous sampling approach The Markup took, they were able to caveat their findings based on biases in their sample. Other research projects (e.g. Siegel et al. (2021); McGrady et al. (2023)) designed identifier-based sampling strategies to create random (to their knowledge) samples of objects on Twitter and YouTube, respectively.
@@
Second, researchers should carefully caveat claims made based on scraped data. There are many factors determining what shows up on a web page or app, particularly when personalization and recommendation systems are in play (e.g., social media feeds). This can mean certain content is over- or under-represented, and that content type may be correlated with unobserved variables (for example, a user’s hobbies and interests). This can introduce a great deal of error in any analysis. At the same time, the content that’s most widely consumed may not be the content that is easily scraped, and when consumption metrics are available they may not be accurate (TikTok, 2023; Herrman, John and Isaac, Mike, 2016). Often, it is not possible to determine what is missing and why, so it may be challenging to statistically de-bias.
@@
Third, researchers should think carefully about how research based on scraped data generalizes. Because researchers often do not have insights into the underlying systems that they are scraping from, it is not often knowable how much the scraped dataset represents (or does not represent) the underlying data, making claims of generalizibility challenging to verify. For example, if a researcher were to scrape search results from Twitter related to a specific set of keywords, there would be no way to verify the completeness of the results or the biases introduced by said incompleteness. Its findings also would not necessarily generalize to other platforms where the audience, affordances, and features may vary. These biases in samples can be particularly difficult to correct using de-biasing approaches since most public information about platform users is highly aggregated.
@@
Rather than tossing out any research where researchers are unable to get a representative sampling frame, researchers should carefully caveat their claims based on the attributes of the dataset they are using. In the other example, where researchers used identifier-based methods to create random samples of users and/or posts, researchers could caveat that their findings are based on an identifier-based sample, and there may be unknown variation in how platforms assign identifiers. Researchers could also consider developing well-defined subsamples (e.g. a group of important political actors, celebrities, news networks, or other groups of interest) and collect full data. This path would typically yield a smaller dataset than a random sample of all posts or all posts with a given set of keywords, but a well-defined sampling frame could reduce uncertainty surrounding unknown biases in the dataset and could offer a clear basis for researchers to make claims about generalizibility (or lack thereof). Because of the limitations of scraping, research based on scraped data should be seen as the first step on a longer journey toward understanding online behavior, with perhaps larger and better samples and/or additional data sources.
@@
While actionable steps will be specific to the particular research project, researchers should consider:
@@
  • A deep assessment of the literature related to said question, particularly in the context of scraping data from platforms in which recommendation systems govern content discovery and distribution
@@ 5 A Researchers’ Checklist
The primary motivation for this paper is to assist researchers in conducting public interest internet research. To that end, while we cannot provide a prescriptive set of recommendations that removes all legal, ethical, institutional, and scientific concerns, we can provide a clear list of considerations for which researchers ought have written answers prior to embarking upon a scraping project. Further, we believe that documenting a researcher’s responses to each of the considerations below will be useful throughout the research process, particularly when engaging with institutional actors as it will help all stakeholders align on the facts and practices of the specific scraping instance. Finally, we recommend that researchers use the information generated in this checklist to generate (with the support of their university, if available) a “communications plan” for how to publicly discuss their activity, particularly in cases where they face backlash for scraping. The full checklist can be found below:
Research Purpose
  • $\square$
    Research question, hypothesis, and other relevant background
  • $\square$
Data Collection
  • $\square$
    Identification of the sources from which the data will originate
  • $\square$
  • $\square$
  • $\square$
  • $\square$
  • $\square$
  • $\square$
    Description of intended collection time frame
Scientific Validity
  • $\square$
  • $\square$
  • $\square$
  • $\square$
Proportionality and Risk Mitigation
  • $\square$
  • $\square$
  • $\square$
  • $\square$
  • $\square$
Stakeholder and Institutional Engagement
  • $\square$
  • $\square$
  • $\square$
@@ 6 Discussion
In the longer term, researchers can and should advocate for protections for researchers using scraped data for public interest research when such research is ethically and scientifically rigorous. For researchers, this can include communicating with policymakers and the public regarding the importance of research on the Internet, as well as demonstrating high ethical rigor to encourage trust in scientists (Dommett and Tromble, 2022). Additionally, researchers can engage with coalitions such as the Coalition for Independent Technology Research [^18] See https://independenttechresearch.org/ to advocate for policies that protect researchers. Researchers should also look for opportunities to engage across the research community to formalize data security and privacy standards for research involving personal information scraped or otherwise obtained from online sources. Data privacy and security standards may ultimately provide a foundation for future data access rights.

### v2

@@ Abstract
Scientists across disciplines often use data from the internet to conduct research, generating valuable insights about human behavior. However, as generative AI relying on massive text corpora becomes increasingly valuable, platforms have greatly restricted access to data through official channels. As a result, researchers will likely engage in more web scraping to collect data, introducing new challenges and concerns for researchers. This paper proposes a comprehensive framework for web scraping in social science research for U.S.-based researchers, examining the legal, ethical, institutional, and scientific factors that we recommend researchers consider when scraping the web. We present an overview of the current regulatory environment impacting when and how researchers can access, collect, store, and share data via scraping. We then provide researchers with recommendations to conduct scraping in a scientifically legitimate and ethical manner. We aim to equip researchers with the relevant information to mitigate risks and maximize the impact of their research amidst this evolving data access landscape.
@@
Legal Disclaimer This paper is intended for informational and academic purposes only. The statements made regarding the law, ethics, and logistics of web scraping are based on general understanding and research as of the publication date. This document does not constitute legal advice, nor does it create an attorney-client relationship. Laws and regulations surrounding web scraping may vary by jurisdiction and are subject to change. Readers are encouraged to consult qualified legal professionals to obtain advice specific to their individual circumstances before engaging in any scraping activities.
@@ 3.1.1 Contractual restrictions
Researchers should exercise caution before concluding that they are not bound by browsewrap agreements that restrict scraping. For starters, in the Berman v Freedom Financial Network, LLC decision, the court noted that if “the website operator can show that a consumer has actual knowledge of the agreement” the terms prohibiting scraping may be enforceable (Berman v. Freedom Financial Network, LLC, 2022, p. 856). What’s more, cases where courts have set aside anti-scraping provisions have typically involved fact patterns where the defendant scraped only publicly-available information or information that the relevant user was already permitted to access through ordinary use of the service. Courts may be less willing to permit scraping that involves content protected by technical safeguards and/or accessible only to certain classes of users. [^5] In this context, researchers’ deploying other users to assist in scraping – such as through the installation of browser extensions that can scrape within their user accounts – could lead to breach of contract claims, at least with regard to those users who have accepted the platform’s user agreement.
@@ 3.4.1 Poor sampling frames
The key takeaway from these sampling frame issues is that the overall representativeness of research based on scraped data is extremely difficult to confirm. For example Tromble et al. (2017) and González-Bailón et al. (2014) find significant variation in Twitter datasets depending on the sampling frame and API endpoints used. With scraped data, however, researchers face a higher level of uncertainty surrounding the sampling frame and data returned, as there is typically no documentation regarding what researchers can expect to reliability collect, as is the case with APIs. Given this limitation, we recommend researchers carefully consider any claims they make with regard to the broad applicability of their findings.
@@ 4 Recommendations
Different data access methods necessitate different legal, scientific, ethical, and institutional considerations. Importantly, the same data collected in different ways may invoke different implications. In the following section, we make recommendations for a set of best practices that U.S.-based researchers can follow (and adapt to their contexts). In addition to immediate recommendations for how to conduct research using scraped data right now, we also provide recommendations for researchers to improve how we take on data scraping projects in the long term.
@@ 4.3 Institutional
As we highlight in Section 3.3, the institutionalization of the ethical, legal, and technical considerations aims to limit risks for researchers, in part by leveraging the expertise that exists across institutional contexts. A key challenge for researchers is determining who to engage, when, and towards what ends. The Institutional Review Board will ultimately be responsible for ensuring that a project conforms with internal policies and legal standards for research. Put simply, we recommend researchers engage their IRB early and often. An IRB should be able to provide guidance on how it understands the risks associated with scraping, as well as connect researchers to other resources from across the university that can help mitigate those risks. Those resources include IT services, libraries, and general counsel. For researchers engaging in scraping for the first time—or for the first time at a particular institution—colleagues can also act as a resource for understanding how scraping is viewed within that particular environment and how to best navigate the specific actors who are involved in institutionalizing the considerations discussed above.
@@ 4.4 Scientific
First, we recommend researchers make as much effort as possible for the project to develop a sampling strategy—as opposed to simply using convenience as their guide. One laudable example is the way in which The Markup developed a sample for their scraping project, Citizen Browser. Unlike many other studies which relied on snowball or other convenience samples for data collection, The Markup started their study by recruiting a panel of Facebook users based on a survey vendor’s national sample. Only then did they install scraping tools connected to those users accounts, which allowed them to appropriately re-weight their data to allow for population inference (The Citizen Browser Project—Auditing the Algorithms of Disinformation – The Markup, 2020). Because of the rigorous sampling approach The Markup took, they were able to caveat their findings based on biases in their sample. Other research projects (e.g. Siegel et al. (2021); McGrady et al. (2023)) designed identifier-based sampling strategies to create random (to their knowledge) samples of objects on Twitter and YouTube, respectively.
@@
Second, we recommend researchers carefully caveat claims made based on scraped data. There are many factors determining what shows up on a web page or app, particularly when personalization and recommendation systems are in play (e.g., social media feeds). This can mean certain content is over- or under-represented, and that content type may be correlated with unobserved variables (for example, a user’s hobbies and interests). This can introduce a great deal of error in any analysis. At the same time, the content that’s most widely consumed may not be the content that is easily scraped, and when consumption metrics are available they may not be accurate (TikTok, 2023; Herrman, John and Isaac, Mike, 2016). Often, it is not possible to determine what is missing and why, so it may be challenging to statistically de-bias.
@@
Third, we recommend researchers think carefully about how research based on scraped data generalizes. Because researchers often do not have insights into the underlying systems that they are scraping from, it is not often knowable how much the scraped dataset represents (or does not represent) the underlying data, making claims of generalizibility challenging to verify. For example, if a researcher were to scrape search results from Twitter related to a specific set of keywords, there would be no way to verify the completeness of the results or the biases introduced by said incompleteness. Its findings also would not necessarily generalize to other platforms where the audience, affordances, and features may vary. These biases in samples can be particularly difficult to correct using de-biasing approaches since most public information about platform users is highly aggregated.
@@
Rather than tossing out any research where researchers are unable to get a representative sampling frame, we recommend researchers carefully caveat their claims based on the attributes of the dataset they are using. In the other example, where researchers used identifier-based methods to create random samples of users and/or posts, researchers could caveat that their findings are based on an identifier-based sample, and there may be unknown variation in how platforms assign identifiers. Researchers could also consider developing well-defined subsamples (e.g. a group of important political actors, celebrities, news networks, or other groups of interest) and collect full data. This path would typically yield a smaller dataset than a random sample of all posts or all posts with a given set of keywords, but a well-defined sampling frame could reduce uncertainty surrounding unknown biases in the dataset and could offer a clear basis for researchers to make claims about generalizibility (or lack thereof). Because of the limitations of scraping, research based on scraped data should be seen as the first step on a longer journey toward understanding online behavior, with perhaps larger and better samples and/or additional data sources.
@@
While actionable steps will be specific to the particular research project, we recommend researchers consider:
@@
  • An assessment of the literature to ensure the sample needed to answer the right set of questions, especially in the context of scraping data from platforms in which recommendation systems govern content discovery and distribution
@@ 5 A Researchers’ Checklist
The primary motivation for this paper is to assist researchers in conducting public interest internet research. To that end, while we cannot provide a prescriptive set of recommendations that removes all legal, ethical, institutional, and scientific concerns, we can provide a clear list of considerations for which we recommend that researchers have written responses prior to embarking upon a scraping project. Further, we believe that documenting a researcher’s responses to each of the considerations below will be useful throughout the research process, particularly when engaging with institutional actors as it will help all stakeholders align on the facts and practices of the specific scraping instance. Finally, we recommend that researchers use the information generated in this checklist to generate (with the support of their university, if available) a “communications plan” for how to publicly discuss their activity, particularly in cases where they face backlash for scraping. We present this checklist for researchers in the supplementary materials.
@@ 6 Discussion
In the longer term, researchers can and should advocate for protections for researchers using scraped data for public interest research when such research is ethically and scientifically rigorous. For researchers, this can include communicating with policymakers and the public regarding the importance of research on the Internet, as well as demonstrating high ethical rigor to encourage trust in scientists (Dommett and Tromble, 2022). Additionally, researchers can engage with coalitions such as the Coalition for Independent Technology Research [^18] See https://independenttechresearch.org/ to advocate for policies that protect researchers. We recommend researchers look for opportunities to engage across the research community to formalize data security and privacy standards for research involving personal information scraped or otherwise obtained from online sources. Data privacy and security standards may ultimately provide a foundation for future data access rights.
@@ Appendix D GDPR
Appendix E Researcher Checklist
Research Purpose
  • $\square$
    Research question, hypothesis, and other relevant background
  • $\square$
Data Collection
  • $\square$
    Identification of the sources from which the data will originate
  • $\square$
  • $\square$
  • $\square$
  • $\square$
  • $\square$
  • $\square$
    Description of intended collection time frame
Scientific Validity
  • $\square$
  • $\square$
  • $\square$
  • $\square$
Proportionality and Risk Mitigation
  • $\square$
  • $\square$
  • $\square$
  • $\square$
  • $\square$
Stakeholder and Institutional Engagement
  • $\square$
  • $\square$
  • $\square$